1.The value of KC for the thermal decomposition of hydrogen sulfide, shown below, is 2.2 × 10-4 at 1400 K.
2H2S(g) ⇌ 2H2(g)+ S2(g)
A sample of gas in which [H2S] = 3.00 M is heated to 1400 K in a sealed vessel. After chemical equilibrium has been achieved, what is the value of [H2S]? Assume no H2 or S2 was present in the original sample. answer in M
1
Expert's answer
2020-04-06T11:48:39-0400

Kc = "{\frac {[H_2]^2[S_2]} {[H_2S]^2}}"

Initial concentration of H2S is 3.00 M. If 2x mol of H2S were involved in reaction, 2x mol of H2 and x mol of S2 were produced. Then in equilibrium state we have:

x mol S2

2x mol H2

3.00 - 2x mol H2S

Kc = "{\frac {[H_2]^2[S_2]} {[H_2S]^2}}={\frac {(2x)^2(x)} {(3-2x)^2}}=2.2*10^{-4}"

"4x^3 = 2.2*10^{-4}*(3-2x)^2=""2.2*10^{-4}*(9+4x^2-12x)=""=1.98*10^{-3}+8.8*10^{-4}x^2-2.64*10^{-3}x"

"4x^3-8.8*10^{-4}x^2+2.64*10^{-3}x-1.98*10^{-3}=0"

After solving the equation (i did it on computer) we got x = 0.0764, then amount of H2S involved in reaction is 2x = 2*0.0764 = 0.1528, in equilibrium [H2S] = 3 - 0.1528 = 2.8472 M